calculate the rate of heat conduction out of the human body, assuming that the core internal temperature is 37.0°C, the skin temperature is 34.0°C, the thickness of the fatty tissues between the core and the skin averages 1.00 cm, and the surface area is 1.40m²


1
Expert's answer
2021-06-16T09:53:03-0400

The thermal conductivity of skin is 0.37 W/(m·K), for fat it is 0.21 W/(m·K), so, the average is 0.29 W/(m·K).

The rate of heat conduction is


"q=\frac{kA(T_2-T_1)}{x}=\frac{0.29·1.40(37-34)}{0.01}=121.8\text{ W}."